1.Determine the reactions at A and B when (a) α = 0, (b) α = 90°, 

(c) α = 30°.


1
Expert's answer
2021-04-12T02:29:38-0400

Consider a simple problem:



The force F is applied at the center of the rod. Hence, the reactions 1 and 2 at A and B will be equal.

(a) α = 0:


"R_1+R_2+F=0,\\
2R+F=0,\\\space\\
R=-\frac F2."

Reactions point to the right (opposite to the force)

(b) α = 90°, 


"R_1+R_2+F\text{sin}90°=0,\\
2R+F=0,\\\space\\
R=-\frac F2."

Reactions point opposite to F downward.

(c) α = 30°


"R_1+R_2+F\text{sin}30°=0,\\
2R+F\text{sin}30°=0,\\\space\\
R=-\frac{F\text{sin}30°}2."

Reactions point to the left bottom corner at the same angle below the horizontal.